OpenStudy (anonymous):

A circle is described by the equation (x−1 2 ) 2 +(y−3 2 ) 2 =4 9 . What are the cordinates for the center of the circle and the length of the radius?

OpenStudy (owlcoffee):

The explicit form of the equation of the circle, has the form: \[(x- \alpha )^{2}+(y- \beta ) ^{2}=r ^{2}\] Where alpha and beta are the coordinates of the center and r is the radius. so if we are given: \[(x-12)^{2}+(y-32)^{2}=49\] We can observe it has the same structure, se we can find the values by saying: \[ \alpha = 12\] \[\beta = 32\] \[r ^{2}=49\]